# Can Contrastive Learning Improve Class-Imbalanced Diffusion Model?

### v1

@@ 4.2 Quantitative Results
      DDPM ddpm       5.81       9.36       7.09       12.64
       +ADA (karras2020training, )       -       -       6.69       12.87
      CBDM (cbdm, )       5.92       9.38       6.52       12.79
      OCLT (oclt, )       5.69       9.42       6.23       13.18
      CLDM (Ours)       5.57       9.42       5.99       13.01
Results Across Various Category Intervals To see the effectiveness of our method on tail classes, we divide the PlacesLT and TinyImageNetLT datasets into three super categories: ‘Head’, ‘Body’, and ‘Tail’, with classes sorted by the number of training images. For each dataset, the top 33% classes were allocated to the ‘Head’ category, the next 34% classes to the ‘Body’ category, and the last 33% classes to the ‘Tail’ category. The percentage $P_{category}$ of training images belonging to each category is shown in the first row of Tab. 3. Our method achieves the best FID score across all three categories on PlacesLT and TinyImageNetLT datasets. For the tail classes, which are only $\sim 3\%$, our method improves their FID from 23.74 to 20.84 on PlacesLT and 40.12 to 37.88 on TinyImageNetLT, respectively. With the proposed MSE loss $\mathcal{L}_{mseclr}$ for positive pairs, our method makes the denoising process insensitive to class conditions for the initial time steps. Thus, tail classes benefit from the rich data of head classes through knowledge sharing, leading to better diversity and fidelity. Additionally, both head and body classes show improved FID scores with our method, as demonstrated in Tab. 3.
Table 3: FID score for three ‘super-categories’: ‘Head’, ‘Body’, and ‘Tail’. All models are measured with DDIM (ddim, ) 100 sampling steps for conditional generation with CFG.
@@
PlacesLT DDPM ddpm 19.26 20.31 23.74 13.89
CBDM cbdm 21.26 21.68 26.07 15.15
OCLT oclt 19.16 19.43 22.84 13.04
CLDM (Ours) 18.15 19.27 20.84 11.99
TinyImageNetLT DDPM ddpm 21.27 33.25 40.12 18.67
CBDM cbdm 24.12 35.02 48.07 20.90
OCLT oclt 20.64 30.62 39.67 17.72
CLDM (Ours) 20.39 31.02 37.88 17.16
@@
Method CIFAR10LT CIFAR100LT
@@
Class-imbalanced Unconditional Generation We also evaluate all the models for class-imbalanced unconditional generation. As shown in Tab. 4, the proposed method reduces FID from 27.52 to 24.16 for CIFAR10LT and from 18.53 to 15.97 for CIFAR100LT. Such an improvement in unconditional generation highlights the effectiveness of the proposed contrastive learning loss, particularly the unsupervised InfoNCE loss $\mathcal{L}_{nceclr}$.
4.3 Qualitative Results
(a) Synthetic images from DDPM (ddpm, ). Red ellipses show similar images generated.
(b) Synthetic images from CLDM (ours) have more diversity and are less repetitive.
Figure 4: The synthesized results for CIFAR100LT tail classes ‘rose’ and ‘table’ are shown for our method and baseline methods. All methods initiate reverse processing from the same Gaussian noise for images at corresponding grid cells. Red dashed ellipses highlight mode collapse issues observed in DDPM. Overall, our method demonstrates higher diversity and fidelity compared to DDPM baseline.
Fig. 4 shows synthetic images of baseline methods and our method for CIFAR100LT tail classes ‘rose’ and ‘table’. All methods start the denoising process from the same Gaussian noise at corresponding grid cells. Red dashed ellipses highlight the mode collapse issues observed in DDPM. With the same initial noise, our method gives synthetic images with higher diversity and fidelity compared to DDPM. For example, DDPM always generates rose images containing only one rose. Our method can generate an image containing two roses (see the sixth column of the last row). To further clarify, in the ‘table’ class, DDPM repeatedly produces near-identical images as highlighted in red ellipses. In contrast, our method produces a more varied set of table images. This suggests that our method can better capture greater diversity in image generation when compared to DDPM (ddpm, ).
Ablation Study
Table 5: Loss Function Analysis on TinyImageNetLT, sampling with DDIM 100 steps.
[…]
$\mathcal{L}_{ddpm}$ 18.67
$\mathcal{L}_{ddpm}+\alpha\cdot\mathcal{L}_{nceclr}$ 18.05
$\mathcal{L}_{ddpm}+\gamma\cdot\mathcal{L}_{mseclr}$ 17.84
$\mathcal{L}_{ddpm}+\alpha\cdot\mathcal{L}_{nceclr}+\gamma\cdot\mathcal{L}_{mseclr}$ 17.16
Table 5 shows an ablation study of the proposed losses including $\mathcal{L}_{nceclr}$ and $\mathcal{L}_{mseclr}$. Our method achieves lower FID when applying only the $\mathcal{L}_{nceclr}$ loss or only $\mathcal{L}_{mseclr}$ loss compared to the original DDPM. The best result is achieved with both losses.
Limitation of our method Our MSE loss involves both conditional generation and unconditional generation, which requires two passes of the denosing network during training and increases training time. In practice, our training time for one epoch is 1.6x of that of DDPM. However, our method doesn’t increase inference latency.
5 Conclusion
Real-world data for training image generation models often exhibit long-tailed distributions. Similar to class-imbalanced GANs (utlo, ), class-imbalanced diffusion models tend to overfit on tail classes due to data scarcity. Inspired by the success and impact of contrastive learning for long-tailed recognition, we are interested in whether contrastive learning can improve the class-imbalanced diffusion model. With well-motivated losses and carefully crafted positive/negative pairs, we effectively leveraged contrastive learning with noticeable improvement. Our framework involves two contrastive losses. Firstly, we employ an unsupervised InfoNCE loss with negative samples only to contrast the latents of different synthetic images at every denoising step, promoting intra-class diversity in particular for tail classes. Our second MSE loss contrasts class-conditional generation with unconditional generation for large timesteps. This encourages the initial denoising steps to be class-agnostic, thereby enriching tail classes through knowledge sharing from head classes–a principle demonstrated to enhance long-tailed GAN performance transitioncgan ; utlo , which we successfully adapt to diffusion models. With the two losses, our contrastive learning method outperforms standard DDPM and other state-of-the-art methods, including CBDM cbdm and OCLT oclt for class-imbalanced diffusion models. Extensive experiments on multiple datasets, including CIFAR10LT, CIFAR100LT, PlacesLT, TinyImageNetLT, and large-scale ImageNetLT, demonstrate the effectiveness and impact of our method.
Appendix A Technical Appendices and Supplementary Material
A.1 Implementation Details
Training Details We keep the same U-Net backbone used in [15] for all methods and strictly follow the training configurations of the baseline models [15, 32, 48], by setting diffusion scheduler $\beta_{1}=10^{-4}$ and $\beta_{T}=0.02$ with T=1000. Adam [49] optimizer is used to optimize the network while the learning rate is set to 0.0002 after 5000 iterations of the warmup process. We use one RTX 4090 GPU to train each model on CIFAR10LT/CIFAR100LT datasets with batch size $64$ while using 2 A100 GPUs for PlacesLT, TinyImageNetLT, and ImageNetLT datasets with batch size $128$. During the training process, all methods are trained from scratch for 300k iterations on ImageNetLT dataset, 200k iterations on CIFAR10LT/CIFAR100LT datasets, 100k iterations on TinyImageNetLT and PlacesLT datasets, while they follow the classifier-free guidance (CFG) algorithm [16], which randomly drops labels with a probability of 10%. For PlacesLT, TinyImageNetLT, and ImageNetLT datasets, we also apply timestep adaptive weight $t/T$ to infoNCE loss since we find the infoNCE loss could also benefit from such an adaptive weight, like MSE loss.
Evaluation Details To evaluate the generation performance for tail classes, we measure FID score for synthetic images of the last 30% classes in each dataset. Specifically, we categorize the ‘Tail’ classes as the last 3 classes for CIFAR10LT, the last 33 classes for CIFAR100LT, the last 66 classes for TinyImageLT (200 classes), the last 121 classes for PlacesLT (365 classes), and the last 333 classes for ImageNetLT (1000 classes), respectively. For CIFAR10LT/CIFAR100LT datasets, the evaluation metrics are based on 50k synthetic images generated by each method, while for the ImageNetLT dataset, the evaluation metrics are based on 20k synthetic images generated by each method. For TinyImageNetLT and PlacesLT dataset, the evaluation metrics are based on 10k synthetic images generated by each method. During inference, we perform a grid search algorithm for each method to determine the optimal guidance strength $\omega$ of CFG, ensuring each model achieves its best performance.
A.2 More Quantitative Results
Table 6: We compare our method with other baselines on CIFAR10LT/CIFAR100LT datasets with DDIM 10 sampling steps for conditional generation with CFG. red ‘()’ shows improvement of our method over DDPM baseline [15]. blue ‘()’ shows improvement of DDPM trained on balanced version over trained on long-tailed version.
[…]
DDPM 14.72 18.27 7.04
CBDM 13.54 16.90 6.52
OCLT 15.48 20.73 7.39
CLDM (Ours) 13.16 (-1.56) 16.83 (-1.44) 6.04 (-1.00)
[…]
DDPM 14.09 25.07 6.13
CBDM 13.37 23.97 5.83
OCLT 13.70 24.48 5.73
CLDM (Ours) 12.90 (-1.19) 23.17 (-1.90) 5.63 (-0.50)
@@ A.2 More Quantitative Results
Table 7: Statistical Analysis for random seed on CIFAR100LT, sampling with DDIM 100 steps.
[…]
@@ A.3 More Qualitative Results
Figure 5: (a) Visualization of low-dimensional embeddings of five training images x for a tail class (‘worm’ in CIFAR100-LT) and synthetic images generated by DDPM [15] and our method. (b) We also show the distributions of real images and synthetic images generated by our method and the original DDPM.
A.3.1 Distribution of Synthetic Images for Tail Class
@@
We visualize synthetic images in the feature space and their density function for our method and the original DDPM. Specifically, we use the Inception-V3 [41] model to extract $2048$ dimensional features of each image, and then apply t-SNE [43] to project these features into $2$ dimensions. As shown in Fig. 5, the original DDPM overfits and generates highly similar images, while synthetic images based on our method have more diversity. We visualize the image distribution more specifically by using kernel density estimation in the bottom. The grey region represents the distribution of real ‘worm’ images from the class-balanced CIFAR100 dataset, while red ‘x’ points denote all 5 ‘worm’ images from the class-imbalanced CIFAR100LT dataset. The red region represents the distribution of images synthesized by our method, while the blue region is for the original DDPM. Areas with higher color saturation (dark, blue, or red) indicate regions of higher density, which correspond to modes of distribution. Synthetic images from DDPM mostly concentrates around the training images, leading to mode collapse. Synthetic images from our method spans the space enclosed by all training images, the distribution of which is shown in red in (b).
Figure 6: Synthesized images of CIFAR10LT tail classes ‘horse,’ ‘ship,’ and ‘truck’ from DDPM. For each column, the images are generated starting from the same Gaussian noise. The synthesized images conditioned on different class labels exhibit similar low-frequency components.
A.3.2 Similarity of Images for Different Classes
We visualize the synthesized images of DDPM for CIFAR10LT tail classes ‘horse’, ‘ship’, and ‘truck’, as shown in Fig. 6. The synthesized images for each column in (a) and (b) are conditioned on different class labels but the same Gaussian noise. It is evident that the synthesized images of each column are coarsely similar and exhibit similar low-frequency components. For instance, in the tenth column of Fig. 6 (a), the synthesized images share a consistent background featuring a purple sky. Similarly, in the fourteenth column (the third column from right to left) of Fig. 6 (b), the images consistently depict an autumn scene with an orange sky. Such observation motivates us to develop knowledge sharing between head and tail classes inspired by the knowledge sharing for long-tailed GAN [22].
Figure 7: To see overfitting on tail classes, we find Top-10 nearest neighbors (3rd to 12th columns sorted by distances) among 1000 synthetic images to an anchor image (Leftmost column) in the training set. KNN is based on $L_{2}$ distances of Inception V3 embeddings. For each example, the top row is the results of DDPM [15], and the bottom row shows ours. The nearest neighbors from DDPM show the mode collapse to the training image, while our method generated more diverse images.
A.3.3 More Visualization of Synthetic Images
@@
To further illustrate how the proposed method mitigates the issue of overfitting, we visualize the top-10 nearest neighbors among 1000 synthetic images to an anchor image in the training set for the original DDPM and our method. As shown in Fig. 7, DDPM shows the mode collapse to the training image while our method’s generated images show much better diversity. For example, in the first two rows, DDPM generates repeated vertical worms while our method generates worms with diverse directions. In the 9th-10th rows, DDPM generates tables with repeated modes while our method generates tables with different styles and colors. We also visualize more synthetic images of CIFAT100LT tail classes for DDPM and our method, as shown in Fig. 8. All these examples illustrate the effectiveness of our method on tail class generation with increased diversity.
(a) DDPM
(b) CLDM (ours)
Figure 8: More synthetic results for CIFAR100LT tail classes. Images in corresponding grid cell for DDPM and our method are initialized from the same Gaussian noise. Our method achieves more diverse images for tail classes.

### v2

@@ Appendix A Technical Appendices and Supplementary Material
DDPM (Ho et al., 2020) 5.93 12.25 1.64
CBDM (Qin et al., 2023) 5.81 10.01 1.58
OCLT (Zhang et al., 2024) 6.10 11.13 1.58
CCUA (ours) 5.56 (-0.37) 10.03 (-2.22) 1.27 (-0.37)
@@
DDPM (Ho et al., 2020) 6.95 17.45 1.71
CBDM (Qin et al., 2023) 6.50 17.36 1.41
OCLT (Zhang et al., 2024) 6.45 17.22 1.42
CCUA (ours) 6.24 (-0.71) 16.35 (-1.10) 1.36 (-0.35)
A.1 Implementation Details
Dataset Details We keep the original $32\times 32$ resolution for CIFAR10-LT/CIFAR100-LT, and resize images to $64\times 64$ for TinyImageNet-LT and Places-LT, while $256\times 256$ for ImageNet-LT. Same as (Qin et al., 2023; Yan et al., 2024; Zhang et al., 2024), we adopt the same methodology presented in (Cao et al., 2019) to construct long-tailed version datasets with an imbalance factor of $0.01$.
Model Architecture Details As described in Sec. 3, the proposed CCUA framework can be applied into UNet-based architecture and Diffusion Transformer. In Fig. 5, we display our UNet-based model and Diffusion Transformer in details. As shown in Fig. 2, the noise estimation network is divided into two parts, a latent encoded network e()e(*) and a decoded network d()d(*). In our setting, for UNet-based model, e()e(*) is defined as the UNet encoder, while d()d(*) is defined as the UNet decoder, as shown in Fig. 5 (a). For Diffusion Transformer, e()e(*) is defined as all the SiT/DiT blocks, while d()d(*) is defined as the final linear and adaLN projection layer, as shown in Fig. 5 (b).
Training Details For SiT pipeline, we strictly follow the same training hyper-parameter settings as Dispersive Loss (Wang & He, 2025). We use 4 A6000 GPUs to train all SiT based models on ImageNet-LT datasets with batch size $48$. During the training process, all methods are trained from scratch. We report 40 epochs, 80 epochs, 120 epochs, and 160 epochs results in Table 1. For DDPM pipeline, we follow the same training configurations of the baseline models (Ho et al., 2020; Qin et al., 2023; Zhang et al., 2024). We use one RTX 4090 GPU to train each model on CIFAR10-LT/CIFAR100-LT datasets with batch size $64$ while using 2 A100 GPUs for TinyImageNet-LT and Places-LT, with batch size $128$. During the training process, all methods are trained from scratch for 200k iterations on CIFAR10-LT/CIFAR100-LT datasets, 100k iterations on TinyImageNet-LT and Places-LT datasets, while they follow the classifier-free guidance (CFG) algorithm (Ho & Salimans, 2022), which randomly drops labels with a probability of 10%. On ImageNet-LT and TinyImageNet-LT datasets, we apply batch resample strategy with the re-balanced factor $0.1$, while we don’t apply batch resample strategy on Places-LT and CIFAR10-LT/CIFAR100-LT datasets. For TinyImageNet-LT and Places-LT datasets, we also apply timestep adaptive weight $t/T$ to unsupervised contrastive loss since we find the unsupervised contrastive loss could also benefit from such an adaptive weight, like alignment loss.
Evaluation Details In Table 2, FIDtail denotes the FID score for synthetic images of the last 30% classes in each dataset. Specifically, we categorize the ‘Tail’ classes as the last 66 classes for TinyImageLT (200 classes), the last 121 classes for Places-LT (365 classes), respectively. In Table 1, the evaluation metrics are based on 50k synthetic images generated by each method with a CFG strength 7.5. In Table 2, the evaluation metrics are based on 10k synthetic images generated by each method. During inference, we perform a grid search algorithm for each method to determine the optimal guidance strength $\omega$ of CFG, ensuring each model achieves its best performance.
A.2 More Quantitative Results
Table 7: We compare our method with other baselines on CIFAR10-LT/CIFAR100-LT datasets, with DDPM 1000 steps for conditional generation with CFG. We also report the data augmentation method ADA (Karras et al., 2020) and $\omega$-scheduler (Wang et al., 2024), which are orthogonal to our method.
   Method    CIFAR10-LT    CIFAR100-LT
[…]
   DDPM (Ho et al., 2020)    5.81    9.36    7.09    12.64
    +ADA (Karras et al., 2020)    -    -    6.69    12.87
    $\omega$-Scheduler (Wang et al., 2024)    5.87    9.22    6.60    12.10
   CBDM (Qin et al., 2023)    5.92    9.38    6.52    12.79
   OCLT (Zhang et al., 2024)    5.69    9.42    6.23    13.18
   CCUA (ours)    5.57    9.42    5.99    13.01
Table 8: We compare our method with other baselines on CIFAR10-LT/CIFAR100-LT datasets with DDIM 10 sampling steps for conditional generation with CFG. Blue ‘()’ shows improvement of our method over DDPM baseline (Ho et al., 2020). Green ‘()’ shows improvement of DDPM trained on balanced version over trained on long-tailed version.
[…]
DDPM 14.72 18.27 7.04
CBDM 13.54 16.90 6.52
OCLT 15.48 20.73 7.39
CCUA (ours) 13.16 (-1.56) 16.83 (-1.44) 6.04 (-1.00)
[…]
DDPM 14.09 25.07 6.13
CBDM 13.37 23.97 5.83
OCLT 13.70 24.48 5.73
CCUA (ours) 12.90 (-1.19) 23.17 (-1.90) 5.63 (-0.50)
[…]
@@ A.2 More Quantitative Results
Method CIFAR10-LT CIFAR100-LT
@@
Class-imbalanced Unconditional Generation We also evaluate all the models for class-imbalanced unconditional generation. As shown in Tab. 9, the proposed method reduces FID from 27.52 to 24.16 for CIFAR10-LT and from 18.53 to 15.97 for CIFAR100-LT. Such an improvement in unconditional generation highlights the effectiveness of the proposed contrastive learning loss, particularly the unsupervised contrastive loss $\mathcal{L}_{ucl}$.
@@
Table 10: Statistical Analysis for random seed on CIFAR100-LT, sampling with DDIM 100 steps.
Seed $\downarrow Method FID\downarrow$
@@
Limitation of our method Our MSE loss involves both conditional generation and unconditional generation, which requires two passes of the denosing network during training and increases training time. In practice, with applying our method, the training time for one epoch is 1.6x of that of DDPM, and 1.48x of that of SiT. However, our method doesn’t increase inference latency.
@@ A.3 More Qualitative Results
A.3.1 More Visualization of Synthetic Images on ImageNet-LT and TinyImageNet-LT
Fig. 6 shows synthetic images of SiT and with the proposed CCUA for ImageNet-LT tail classes ‘bubble’, ‘redwine’, ‘comic book’ and ‘yawl’. Fig. 7 shows synthetic images of DDPM and with the proposed CCUA for TinyImageNet-LT tail classes ‘teapot’, ‘water tower’, ‘pretzel’, ‘mushroom’, ‘orange’, and ‘pizza’. These methods start the denoising process from the same Gaussian noise at corresponding grid cells. As shown in Fig. 6 and Fig. 7, synthetic images of CCUA show consistently higher diversity and fidelity compared to SiT.
(a) Synthetic images from SiT (Ma et al., 2024).
(b) Synthetic images from CCUA (ours).
Figure 6: Synthetic images of SiT and CCUA for ImageNet-LT tail classes (from top-left to right-bottom: ‘bubble’, ‘redwine’, ‘comic book’ and ‘yawl’). All methods start the denoising process from the same Gaussian noise at corresponding grid cells. CCUA shows consistently higher diversity and fidelity compared to SiT.
(a) Synthetic images from DDPM (Ho et al., 2020).
(b) Synthetic images from CCUA (ours).
Figure 7: More synthetic results for TinyImageNet-LT tail classes (from top-left to right-bottom: ‘teapot’, ‘water tower’, ‘pretzel’, ‘mushroom’, ‘orange’, and ‘pizza’). Images in corresponding grid cell for DDPM and our method are initialized from the same Gaussian noise. Our method achieves more diverse images with higher fidelity for tail classes. Note that for ‘pretzel’ and ‘orange’ classes, DDPM fails to synthesize images correlated to the class while CCUA synthesizes diverse images with high quality.
@@ A.3.1 More Visualization of Synthetic Images on ImageNet-LT and TinyImageNet-LT
Figure 8: (a) Visualization of low-dimensional embeddings of five training images x for a tail class (‘worm’ in CIFAR100-LT) and synthetic images generated by DDPM (Ho et al., 2020) and our method. (b) We also show the distributions of real images and synthetic images generated by our method and the original DDPM.
A.3.2 Distribution of Synthetic Images for Tail Class
@@
We visualize synthetic images in the feature space and their density function for our method and the original DDPM. Specifically, we use the Inception-V3 (Szegedy et al., 2016) model to extract $2048$ dimensional features of each image, and then apply t-SNE (Van der Maaten & Hinton, 2008) to project these features into $2$ dimensions. As shown in Fig. 8, the original DDPM overfits and generates highly similar images, while synthetic images based on our method have more diversity. We visualize the image distribution more specifically by using kernel density estimation in the bottom. The grey region represents the distribution of real ‘worm’ images from the class-balanced CIFAR100 dataset, while red ‘x’ points denote all 5 ‘worm’ images from the class-imbalanced CIFAR100-LT dataset. The blue region represents the distribution of images synthesized by our method, while the green region is for the original DDPM. Areas with higher color saturation (dark, green, or blue) indicate regions of higher density, which correspond to modes of distribution. Synthetic images from DDPM mostly concentrates around the training images, leading to mode collapse. Synthetic images from our method spans the space enclosed by all training images, the distribution of which is shown in blue in (b).
A.3.3 Mode Collapse Issue on Tail Class
@@
Fig. 9 shows synthetic images of baseline methods and our method for CIFAR100-LT tail classes ‘rose’ and ‘table’. All methods start the denoising process from the same Gaussian noise at corresponding grid cells. Red dashed ellipses highlight the mode collapse issues observed in DDPM. With the same initial noise, our method gives synthetic images with higher diversity and fidelity compared to DDPM. For example, DDPM always generates rose images containing only one rose. Our method can generate an image containing two roses (see the sixth column of the last row). To further clarify, in the ‘table’ class, DDPM repeatedly produces near-identical images as highlighted in red ellipses. In contrast, our method produces a more varied set of table images. This suggests that our method can better capture greater diversity in image generation when compared to DDPM (Ho et al., 2020). To further illustrate how the proposed method mitigates the issue of overfitting, we visualize the top-10 nearest neighbors among 1000 synthetic images to an anchor image in the training set for the original DDPM and our method. As shown in Fig. 10, DDPM shows the mode collapse to the training image while our method’s generated images show much better diversity. For example, in the first two rows, DDPM generates repeated vertical worms while our method generates worms with diverse directions. In the 9th-10th rows, DDPM generates tables with repeated modes while our method generates tables with different styles and colors.
(a) Synthetic images from DDPM (Ho et al., 2020). Red ellipses show similar images generated.
(b) Synthetic images from CCUA (ours) have more diversity and are less repetitive.
Figure 9: The synthesized results for CIFAR100-LT tail classes ‘rose’ and ‘table’ are shown for our method and baseline methods. All methods initiate reverse processing from the same Gaussian noise for images at corresponding grid cells. Red dashed ellipses highlight mode collapse issues observed in DDPM. Overall, our method demonstrates higher diversity and fidelity compared to DDPM baseline.
Figure 10: To see overfitting on tail classes, we find Top-10 nearest neighbors (3rd to 12th columns sorted by distances) among 1000 synthetic images to an anchor image (Leftmost column) in the training set. KNN is based on $L_{2}$ distances of Inception V3 embeddings. For each example, the top row is the results of DDPM, and the bottom row shows ours. The nearest neighbors from DDPM show the mode collapse to the training image, while our method generated more diverse images.